• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:09:48 on November 7,2002

    Entry number 89806

    keyword=Shift Summary

    Day Shift:
    John LeRose
    Patricia Solvignon
    George Chang

    Production running at Q3D_h20
    Left Arm: electrons, P0=3.5635GeV/c, 25.588 degrees
    Right: protons, P0=2.1627GeV/c, 40.960 degrees
    Run 3054 in progress at start of shift
    8:57 run 3055 started
    9:49 run 3056 started
    10:08 MCC taking beam for beam studies (ended run 3056)
    Still need to do optics target work on Q3D_h20.
    13:55 optics target inserted
    14:15 3057 started optics run (prescales set to 2,1,1,2,big)
    14:25 LD2 target inserted
    14:30 3058 started (efficiency run prescalesT1=T2=T3=1)
    14:37 going to kinematics Q3D_hiN
    Left arm unchanged
    right arm 49.00 degrees P=2.0285 GeV/c
    15:09 run 3059 (Junk prescale file not saved)
    15:15 run 3060 (ps= 1/3/3/10/6553..)
    15:30 ended 3060 going to kinematics Q3D_i40
    Right unchanged
    Left 25.373 degrees P0= 2.6254
    Problems submitting beam time use with new bta (said I had errors that weren't there. Requested that I correct them and resubmit but then deleted everything automatically and setup for the next shift)
    Fortunately, I kept a parallel record with the old setup and submitted that successfully.
    At run end swing shift finishing prep for newkinematics